OpenStudy (adamsaleh):

prove the following: 1. (1-sinxtanx)/1+secx=cosx 2. (1+sinx)/cosx =cosx/1-sinx

OpenStudy (ikram002p):

\(\Huge \frac {(1-\sin x\tan x)}{1+\sec x}=\cos x \)

OpenStudy (anonymous):

LHS IS NOT EQUAL TO RHS

OpenStudy (anonymous):

The two sides are not equal

OpenStudy (ikram002p):

ur correct No.name

OpenStudy (anonymous):

In the second problem Multiply both the num and denom by (1-sinx) In the numerator it would become cos^2 x as (1+sinx)(1-sinx) = cos^2 x Which you can write as cos x cos x On the RHS multiply by the numerator by cos x and cos x You would get cos x cos x /cos x (1 - sin x ) Which is the same as LHS hence proved